Circular No. 8939

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 G2 (SHOEMAKER) Comet P/1994 J3 = 1994k = 1994 XXVIII (Shoemaker, a.k.a. ‘Shoe-

maker 4’; cf. IAUC 5991, 5998, etc.) was identified by T. B. Spahr in single-night observations of minor planets (first observation tabulated below) that were reported by the Catalina Sky Survey to the Minor Planet Center, with additional observations found by B. G. Marsden (see CBET 1347 for additional details). The indicated correction to the prediction on MPC 56803 (ephemeris on MPC 60734) is ΔT = −1.7 days. The available astrometry, orbital elements, and ephemeris appear on MPEC 2008-H20.

2008 UT  α_{2000}  δ_{2000}  Mag.
Apr. 10.30664  15^{h}06^{m}27.^{s}21  +11^{o}53′06′′.5  18.6

SUPERNOVA 2008bo IN NGC 6643 C. J. Stockdale, Marquette University and University of Oklahoma;

K. W. Weiler, Naval Research Laboratory; S. Immler, Goddard Space Flight Center; J. M. Marcaide, University of Valencia; N. Panagia, Space Telescope Science Institute and Observatory of Catania; S. D. Van Dyk, Spitzer Science Center, California Institute of Technology; R. A. Sramek, National Radio Astronomy Observatory; and D. Pooley, University of Wisconsin at Madison, report the detection of radio emission from the type-II supernova 2008bo (cf. CBETs 1324, 1325) with the Very Large Array (VLA) radio telescope. The measured position of the radio emission of α = 18^{h}19^{m}54^{s}.41, δ = +74^{o}34′21′′.0, (equinox 2000.0) is in good agreement with the measured optical position (end figures 54^{s}.34, 20′′.9, from CBET 1324) and with the previously reported radio position (end figures 54^{s}.30,

21′′.1, from http://www.astronomerstelegram.org/?read=1477). A flux

density of 0.507 ± 0.043 mJy was measured at 8.46 GHz (wavelength 3.5 cm) on Apr. 18.52 UT. This observation supports the first radio detection (op.cit.), which had a flux density of 0.234 ± 0.070 mJy measured at 22.46 GHz (wavelength 1.3 cm) and 0.477 ± 0.053 mJy at 8.46 GHz (wavelength 3.5 cm) on Apr. 5.75. No radio emission was detected near the vicinity of SN 2008bo with a 3σ upper limit of 0.23 mJy at 8.46 GHz on Apr. 7.10. The recent radio measurements presented here were prompted by the x-ray brightening (ATEL 1481, 1463) detected in the Swift observations of SN 2008bo. This observed radio re-brightening is very uncommon. Further radio observations have been requested. VLA scheduling officers generously supported the authors’ observing program.

2008 April 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green